OpenStudy (anonymous):

What is the equation of the hyperbola with a=6 and c=10. Assume that the transverse axis is horizonatial

OpenStudy (amistre64):

solve a pythag for b and plug it all in

OpenStudy (anonymous):

I am lost

OpenStudy (amistre64):

on what?

OpenStudy (anonymous):



OpenStudy (amistre64):

theres something i can work with :)

OpenStudy (amistre64):

the 100 is wrong ... and you need and =1 tacked onthe end

OpenStudy (amistre64):

a hyperbola does a normal looking pythag reference (its the ellipse thats wierd on it) \[a^2+b^2 = c^2\] \[36+b^2 = 100\] \[b^2 = 100-36\]

OpenStudy (anonymous):

ok so 6^2+64^2=10^2?

OpenStudy (amistre64):

b=8, yes therefore under y is 64

OpenStudy (amistre64):